# Supporting SENĆOT EN Language Documentation Efforts with Automatic Speech Recognition

### v1

@@ 5.3 Performance Analysis
1 | Wav2Vec2-random-int | - | - | 84.36 | 99.94 | 100.00 | 99.96 |
2 | Wav2Vec2-base | ✗ | - | 10.68 | 36.18 | 65.99 | 49.10 |
3 | Wav2Vec2-large | 8.25 | 21.42 | 56.87 | 35.04 | ||
4 | Wav2Vec2-xlsr-53 | ✓ | - | 11.23 | 33.48 | 69.13 | 51.44 |
5 | Wav2Vec2-xls-r-300m | 10.41 | 31.55 | 63.35 | 45.63 | ||
6 | Wav2Vec2-xls-r-1b | 6.32 | 14.87 | 55.04 | 27.81 | ||
7 | Data2Vec-base | ✗ | - | 14.89 | 40.09 | 78.56 | 60.61 |
8 | Data2Vec-large | 9.29 | 27.75 | 60.08 | 40.51 | ||
9 | HuBERT-base | ✗ | - | 13.34 | 45.15 | 72.04 | 58.61 |
10 | HuBERT-large | 12.29 | 44.66 | 69.78 | 56.06 | ||
11 | WavLM-base | ✗ | - | 11.70 | 36.18 | 71.27 | 50.71 |
12 | WavLM-large | 13.43 | 45.98 | 71.88 | 59.61 | ||
13 | Whisper-medium-en | ✗ | - | 7.36 | 15.38 | 57.30 | 28.13 |
14 | Whisper-large-v2 | ✓ | 7.11 | 14.60 | 58.01 | 27.66 | |
15 | Wav2Vec2-xls-r-1b | ✓ | small-4g | 6.05 | 12.18 | 57.92 | 25.13 |
16 | large-4g | 5.63 | 12.35 | 49.09 | 23.16 | ||
17 | Whisper-large-v2 | ✓ | small-4g | 6.53 | 11.09 | 60.14 | 25.16 |
18 | large-4g | 6.12 | 10.95 | 50.35 | 22.67 |
@@ Appendix A Further Ablation Studies
1 | ✗ | 1.7 | small-4g | 19.68 | 18.93 | 100.00 | 46.92 |
2 | large-4g | 19.59 | 20.46 | 100.00 | 49.34 | ||
3 | ✓ | 13.3 | small-4g | 19.72 | 16.67 | 100.00 | 46.23 |
4 | large-4g | 9.65 | 16.62 | 58.92 | 36.63 |

### v2

@@ 5.3 Performance Analysis
| 1 | Wav2Vec2-random-int | - | - | 84.36 | 99.94 | 100.00 | 99.96 |
| 2 | Wav2Vec2-base | ✗ | - | 10.68 | 36.18 | 65.99 | 49.10 |
| 3 | Wav2Vec2-large | 8.25 | 21.42 | 56.87 | 35.04 | ||
| 4 | Wav2Vec2-xlsr-53 | ✓ | - | 11.23 | 33.48 | 69.13 | 51.44 |
| 5 | Wav2Vec2-xls-r-300m | 10.41 | 31.55 | 63.35 | 45.63 | ||
| 6 | Wav2Vec2-xls-r-1b | 6.32 | 14.87 | 55.04 | 27.81 | ||
| 7 | Data2Vec-base | ✗ | - | 14.89 | 40.09 | 78.56 | 60.61 |
| 8 | Data2Vec-large | 9.29 | 27.75 | 60.08 | 40.51 | ||
| 9 | HuBERT-base | ✗ | - | 13.34 | 45.15 | 72.04 | 58.61 |
| 10 | HuBERT-large | 12.29 | 44.66 | 69.78 | 56.06 | ||
| 11 | WavLM-base | ✗ | - | 11.70 | 36.18 | 71.27 | 50.71 |
| 12 | WavLM-large | 13.43 | 45.98 | 71.88 | 59.61 | ||
| 13 | Whisper-medium-en | ✗ | - | 7.36 | 15.38 | 57.30 | 28.13 |
| 14 | Whisper-large-v2 | ✓ | 7.11 | 14.60 | 58.01 | 27.66 | |
| 15 | Wav2Vec2-xls-r-1b | ✓ | small-4g | 6.05 | 12.18 | 57.92 | 25.13 |
| 16 | large-4g | 5.63 | 12.35 | 49.09 | 23.16 | ||
| 17 | Whisper-large-v2 | ✓ | small-4g | 6.53 | 11.09 | 60.14 | 25.16 |
| 18 | large-4g | 6.12 | 10.95 | 50.35 | 22.67 | ||
@@ Appendix A Further Ablation Studies
| 1 | ✗ | 1.7 | small-4g | 19.68 | 18.93 | 100.00 | 46.92 |
| 2 | large-4g | 19.59 | 20.46 | 100.00 | 49.34 | ||
| 3 | ✓ | 13.3 | small-4g | 19.72 | 16.67 | 100.00 | 46.23 |
| 4 | large-4g | 9.65 | 16.62 | 58.92 | 36.63 | ||